OpenStudy (anonymous):

Solve the system. 2x+3y-2z=-5 -2x+2y+3z=6 x+y+z=2

OpenStudy (campbell_st):

easiest thing to do is eliminate x add the 1st and 2nd equations and you get 2x + 3y - 2z = -5 -2x +2y +3z = 6 --------------------- 5y + z = 1 equation 4 take the 3rd equation multiply it by -2 and add equation 1 2x + 3y -2z = -5 -2x -2y -2z = -4 ---------------- y -4z = -9 Equation 5 make y the subject in equation 5 and substitute it into equation 4 y = -9 + 4z so substituting 5(-9 + 4z) + z = 1 -45 + 20z + z = 1 just solve for z then substitute into equation 4 to find y then substitute into equation 1 to find x all the answers involve fractions...

OpenStudy (anonymous):

Ok. I got those. I didn't think I would have fractions. Thanks.

OpenStudy (anonymous):

(1/21, -5/21, 46/21)